# Bandits with Abstention under Expert Advice

### v1

@@ 6 Efficient Learning with Metric Balls
We now give a simple example: that of finding disjoint balls in a finite space endowed with a distance function. We will also give subroutines that speed up CBA in this problem. Specifically, we have a set of contexts $\mathcal{X}$ with $|\mathcal{X}|=N$ for some $N\in\mathbb{N}$ , as well as a distance function $d:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$ known to the learner a-priori. A-priori nature selects a sequence: $\langle(x_{t},\boldsymbol{r}_{t})\in\mathcal{X}\times[-1,1]^{K}\,|\,t\in[T]\rangle$, but does not reveal it to the learner. For all $t\in[T]$ we define $r_{t,\square}:=0$. On each trial $t\in[T]$ the following happens:
  1. 1.
    The context $x_{t}$ is revealed to the learner.
  2. 2.
    the learner selects an action $a_{t}\in[K]\cup\{\square\}$.
  3. 3.
    the learner sees and incurs reward $r_{t,a_{t}}\in[-1,1]$.
Our cumulative reward bound will be based on the notion of a ball. Specifically, a ball is any set $\mathcal{B}\subseteq\mathcal{X}$ in which there exists some $x\in\mathcal{X}$ and $\delta\in\mathbb{R}_{+}$ with:
=$\mathcal{B}=\{z\in\mathcal{X}\,|\,d(x,z)\leq\delta\}.$
We now give the reward and time complexity bounds for the implementation of CBA that will follow:
Theorem 6.1.
Given any $M\in\mathbb{N}$ we can implement CBA such that for any sequence of disjoint balls $\langle\mathcal{B}_{j}\,|\,j\in[M]\rangle$ with corresponding actions $\langle b_{j}\in[K]\,|\,j\in[M]\rangle$ we have:
$\sum_{t\in[T]}\mathbb{E}[r_{t,a_{t}}]\geq$ $\sum_{t\in[T]}\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket r_{t,b_{j}}$
$-\sqrt{4M\ln(N)(6K+1)T}.$
We briefly comment on the term: $\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket r_{t,b_{j}},$ that appears in the theorem statement. If $x_{t}$ does not belong to any of the balls in $\langle\mathcal{B}_{j}\,|\,j\in[M]\rangle$ then this term is equal to zero (which is the reward of abstaining). Otherwise, since the balls are disjoint, $x_{t}$ belongs to exactly one of them and the term is equal to the reward induced by the action that corresponds to that ball. In other words, the total cumulative reward is bounded relative to that of the policy that abstains whenever $x_{t}$ is outside the union of the balls and otherwise selects the action corresponding to the ball that $x_{t}$ lies in.
Note that the vast improvement in our reward bound is over that of SpecialistExp with abstention as one of the actions. In order to get a reward bound for SpecialistExp, the sets in which the specialists are awake must partition the set $\mathcal{X}$. This means that we must add to our $M$ balls a disjoint covering (by balls) of the complement of the union of the original $M$ balls. Remark that the added balls correspond to the sets in which the specialists predicting the abstention action are awake. Typically this would require a huge number of balls so that the total number of specialists is huge (much larger than $M$); this huge number of specialists essentially replacing the term $M$ in our reward bound. Furthermore, in Appendix A, we show that the same implementation of CBA is capable of learning a weighted set of overlapping balls, as long as the sum of the weights of the balls covering any context is bounded above by one, which SpecialistExp cannot do in general.
The choice of experts for CBA that leads to Theorem 6.1 is defined by the set of ball-action pairs so that $E=N^{2}K$ and for each ball $\mathcal{B}$ and action $a\in[K]$ there exists an unique $i\in[E]$ such that for all $t\in[T]$ and $b\in[K]$ we have: $e^{i}_{t,b}:=\llbracket x_{t}\in\mathcal{B}\rrbracket\llbracket b=a\rrbracket\,.$ By choosing $w_{1,i}=M/N^{2}$ for all $i\in[E]$ , and choosing $\eta:=(M\ln(N)/(6K+1)T)^{-1/2}\,.$ Theorem 3.1 implies the reward bound in Theorem 6.1.
The per-trial time complexity of a direct implementation of CBA for this set of experts would be $\mathcal{O}(KN^{2})$. We now show how to implement CBA in a per-trial time of only $\mathcal{O}(KN\ln(N))$. To do this first note that we can assume, without loss of generality, that for all $q,x,z\in\mathcal{X}$ with $x\neq z$ we have $d(q,x)\neq d(q,z)$ since ties can be broken arbitrarily and balls can be duplicated.
For all $x,z\in\mathcal{X}$ and $a\in[K]$ we have the following:
For all $a\in[K]$ we have:
Finally, for all $x,z\in\mathcal{X}$ and $a\in[K]$ we have the following:
[…]
@@
(a) Stochastic Block Model
(b) Gaussian graph
(c) Cora graph
(d) LastFM Asia graph
Figure 1: Results regarding the number of mistakes over time, the three main settings are presented from left to right: the Stochastic Block Model, Gaussian graph, Cora graph and LastFM Asia graph. In this context, D1, D2, and D-INF represent the $p$-norm bases, LVC represents the community detection basis, and INT represents the interval basis. The baselines, EXP3 for each context, Contextual Bandit with similarity, and GABA-II, are denoted as EXP3, CBSim, and GABA, respectively, and are represented with dashed lines. All the figures, display the data with 95% confidence intervals over 20 runs, calculated using the standard error multiplied by the $z$-score 1.96.
7 Preliminary Experiments
We empirically evaluate our approach in the context of online multi-class node classification on a given graph with bandit feedback. At each time step, the algorithm is presented with a node chosen uniformly at random and must either predict an action from the set of possible actions $[K]$ or abstain. The node can accept (resulting in a positive reward) or reject (resulting in a negative reward) the suggestion based on its preferred class with a certain probability. This can be seen as a special case of the metric space example from Section 6, where the context set $\mathcal{X}=[N]$ corresponds to the node set. In general, we are not constrained to metric balls but can use arbitrary experts that correspond to set-action pairs of the form $(\mathcal{B},k)\in 2^{\mathcal{X}}\times[K]$. We call any such set-action pair a basis set, and a set of basis sets a basis. For our experiments, we selected some representative bases used on learning tasks on graphs before, each leading to different inductive priors on the node labeling. For this reason, we will distinguish the different classes as foreground classes (classes in which the learning bias structure is supposed to hold) and background classes (nodes for which is better to abstain). We provide a short description of the basis sets here and refer to Appendix B for more details.
Effective $p$-resistance basis $d_{p}$: Balls given by the metric
We use $d_{1}$, $d_{2}$, and $d_{\infty}$ (Herbster & Lever, 2009).
Louvain method basis (LVC): Communities returned by the Louvain method (Blondel et al., 2008), processed by the greedy peeling algorithm (Lanciano et al., 2023).
Geodesic intervals basis (INT): All sets of the form $I(x,y):=\{z\in\mathcal{X}\mid z\text{ is on a shortest }x\text{-}y\text{ path}\}$ for all $x,y\in\mathcal{X}$ (Pelayo, 2013; Thiessen & Gärtner, 2021).
Stochastic Block Model. This graph, inspired by Holland et al. (1983), is generated by spawning an arbitrary number of disjoint cliques representing the foreground classes. Then an arbitrary number of background points are generated and connected to every possible point with a low probability. In Figure 1(a) are displayed the results for the case of $F=160$ nodes for each foreground class and $B=480$ nodes for the background class. Connecting each node of the background class with a probability of $1/\sqrt{FB}$.
Real-world dataset. We tested our approach on the Cora dataset (Sen et al., 2008) and the LastFM Asia dataset (Leskovec & Krevl, 2014). While both of these graphs contain both features and a graph, we exclusively utilized the largest connected component of each graph, resulting in 2485 nodes and 5069 edges for the Cora graph and 7624 nodes and 27806 edges for the LastFM Asia graph. Subsequently, we randomly chose a subset of three out of the original seven and eighteen classes, respectively, to serve as the background class. Additionally, we selected 15% of the nodes from the foreground classes randomly to represent noise points, and we averaged the results over multiple runs, varying the labels chosen for noise. Both in Figures 1(c) and 1(d) we averaged over 5 different label sets as noise. For the LastFM Asia graph, we exclusively tested the LVC bases, as it is the most efficient one to compute given the large size of the graph.
Results. The results from both synthetically generated tests (Figures 1(a) and 1(b)) demonstrate the superiority of our method when compared to the baselines. In particular, $d_{\infty}$-balls delivered exceptional results for both graphs, implying that $d_{\infty}$-balls effectively cover the foreground classes as expected. For the Cora dataset (Figure 1(c)), we observed that our method outperforms GABA-II only when employing the community detection basis set. This similarity in performance is likely attributed to the dataset’s inherent lack of noise. Worth noting that the method we employed to inject noise into the dataset may not have been the optimal choice for this specific context. However, it is essential to highlight that our primary focus revolves around the abstention criteria, which plays a central role in ensuring the robustness of our model in the presence of noise. For the LastFM Asia dataset, our objective was to assess the practical feasibility of the model on a larger graph. We tested the LVC bases as they were the most promising and most efficient to compute. We outperform the baselines in our evaluation as shown in Figure 1(d) and further discussed in Appendix C.
In summary, our first results confirm what we expected: our approach excels when we choose basis functions that closely match the context’s structure. However, it also encounters difficulties when the chosen basis functions are not a good fit for the context. In Appendix C, the results for a wide range of different parameters used to generate the previously described graphs are displayed.
Impact Statement
Appendix A Overlapping balls extension
@@
In this section, we present the theorem that allows us to present the results of overlapping balls as expressed in Section 6. Note that Theorem 6.1 is the special case of Theorem A.1 when the balls are disjoint and $u_{j}=1$ for all $j\in[M]$.
Theorem A.1.
@@
Let $M\in\mathbb{N}$ and $\{(\mathcal{B}_{j},b_{j},u_{j})\,|\,j\in[M]\}$ be any sequence such that $\mathcal{B}_{j}$ is a ball, $b_{j}\in[K]$ is an action, and $u_{j}\in[0,1]$ is such that for all $x\in\mathcal{X}$ we have:
$\sum_{j\in[M]}\llbracket x\in\mathcal{B}_{j}\rrbracket u_{j}\leq 1\,.$
@@
${r}^{*}_{t}:=\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket u_{j}{r}_{t,b_{j}}\,,$
which represents the reward of the policy induced by $\{(\mathcal{B}_{j},b_{j},u_{j})\,|\,j\in[M]\}$ on trial $t$. The regret of CBA, with the set of experts given in Section 6 and with correctly tuned parameters, is then bounded by:
@@
Direct from Theorem 3.1 using the experts (with efficient implementation) given in Section 6
Appendix B The Details of the Graph Bases
This section expands the definition and explanations for the bases we used in the Experiment. Remember that we refer to any set of experts that correspond to set-action pairs of the form $(\mathcal{B},k)\in 2^{\mathcal{X}}\times[K]$ as a basis set, and a set of basis sets as basis.
B.1 $p$-seminorm Balls on Graphs
As we see in Sec. 6, the CBA seems to work only for vector data. However, in the following sections, we explore how our CBA algorithm can be applied to graph data by creating a ball structure over the graph.
@@ B.1 $p$-seminorm Balls on Graphs
which is a well-defined metric for $p\in[1,\infty)$ if the graph is connected and may be defined for $p=\infty$ by taking the appropriate limits. When $p=2$ this is the square root of the effective resistance circuit between nodes $i$ and $j$ which comes from interpreting the graph as an electric circuit where the edges are unit resistors and the denominator of Equation (16) is the power required to maintain a unit voltage difference between $u$ and $v$ (Doyle & Snell, 1984). More generally, $d_{p}(i,j)^{p}$ is known as $p$-(effective) resistance (Herbster & Lever, 2009; Alamgir & von Luxburg, 2011; Saito & Herbster, 2023). When $p\in\{1,2,\infty\}$ there are natural interpretation of the $p$-resistance. In the case of $p=1$, we have that the effective is equal to one over the number of edge-disjoint paths between $i$ and $j$ which is equivalently one over the minimal cut that separates $i$ from $j$. When $p=2$ it is the effective resistance as discussed above. And finally when $p=\infty$ we have that $d_{\infty}$ is the geodesic distance (shortest path) between $i$ and $j$. Note that, interestingly, there are at most $2N$ distinct balls for $d_{1}$; as opposed to the general bound $O(N^{2})$ on the number of metric balls. This follows since $d_{1}$ is an ultrametric. A nice feature of metric balls is that they are ordinal, i.e., we can take an increasing function of the distance and the distinct are unchanged. The time complexity for each ball is as follows. For $d_{1}$ ball, we compute every pair of distance in $\mathcal{O}(N^{3})$ using the Gomory-Hu tree (Gomory & Hu, 1961). For $d_{2}$ ball, it is actually enough to compute the pseudoinverse of graph Laplacian once, which costs $\mathcal{O}(N^{3})$ (Doyle & Snell, 1984). For $d_{\infty}$ ball, we can compute every pair of distance in $\mathcal{O}(N^{3})$ by Floyd–Warshall algorithm (Floyd, 1962).
B.2 Community detection bases
@@
We can compute a basis for a given graph $G=(V,E)$ using community detection algorithms. Community detection is one of the most well-studied operations for graphs, where the goal is to find a partition $\{C_{1},\dots,C_{q}\}$ of $V$ (i.e., $\bigcup_{i=1}^{q}C_{i}=V$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$) so that each $C_{i}$ is densely connected internally but sparsely connected to the rest of the graph (Fortunato, 2010). There are many community detection algorithms, all of which can be used here, but the most popular algorithm is the Louvain method (Blondel et al., 2008). We briefly describe how this algorithm works. The algorithm starts with an initial partition $\{\{v\}\mid v\in V\}$ and aggregates the clusters iteratively: For each $v\in V$, compute the gain when moving $v$ from its current cluster to its neighbors’ clusters and indeed move it to a cluster with the maximum gain (if the gain is positive). Note that the gain is evaluated using modularity, i.e., the most popular quality function for community detection (Newman & Girvan, 2004). The algorithm repeats this process until no movement is possible. Then the algorithm aggregates each cluster to a single super node (with appropriate addition of self-loops and change of edge weights) and repeats the above process on the coarse graph as long as the coarse graph is updated. Finally, the algorithm outputs the partition of $V$ in which each cluster corresponds to each super node in the latest coarse graph. Note that it is widely recognized that the Louvain method works in $\mathcal{O}(N\log N)$ in practice (Traag, 2015).
@@
To obtain a finer-grained basis set, we apply the so-called greedy peeling algorithm for each $C_{i}$ in the output of the Louvain method. For $C_{i}\subseteq V$ and $v\in C_{i}$, we denote by $d_{C_{i}}(v)$ the degree of $v$ in the induced subgraph $G[C_{i}]$. For $G[C_{i}]$, the greedy peeling iteratively removes a node with the smallest degree in the currently remaining graph and obtains a sequence of node subsets from $C_{i}$ to a singleton. Specifically, it works as follows: Set $j\leftarrow|C_{i}|$ and $C_{i}^{(j)}\leftarrow C_{i}$. For each $j=|C_{i}|,\dots,2$, compute $v_{\text{min}}\in\operatorname*{arg\,min}\{d_{C_{i}^{(j)}}(v)\mid v\in C_{i}^{(j)}\}$ and $C_{i}^{(j-1)}\leftarrow C_{i}^{(j)}\setminus\{v_{\text{min}}\}$. Using a sophisticated data structure, this algorithm runs in linear time (Lanciano et al., 2023).
@@
In summary, our community detection basis set is the collection of node subsets $\{C_{i}^{(j)}\mid i=1,\dots,q,\,j=1,\dots,|C_{i}|\}$ together with $\{\{v\}\mid v\in V\}$ for completeness.
B.3 Graph convexity bases

### v2

@@ Appendix B Efficient implementation proof
Appendix C Lower bound proof
Proposition C.1.
Take any learning algorithm. Given any basis $\mathcal{B}$ and any $M\in\mathbb{N}$ then for any sequence of disjoint basis elements $\langle B_{j}\,|\,j\in[M]\rangle$ there exists a sequence of corresponding actions $\langle b_{j}\in[K]\,|\,j\in[M]\rangle$ such that an adversary can force:
$\sum_{t\in[T]}\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket r_{t,b_{j}}-\sum_{t\in[T]}\mathbb{E}[r_{t,a_{t}}]\in\Omega(\sqrt{MKT})$
Proof.
In this scenario, at each time step, either a single expert (i.e., the basis element containing the current context $x_{t}$) is active, making predictions based on its label, or no expert is active, prompting the learner to abstain and thus incur zero reward or cost.
Therefore we define $T^{\prime}=\{t\in[T]\,|\,\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket=1\}$ as the set of timesteps in which the learner is going to play. Since the concept of abstention is that our algorithm is not going to pay anything for the timesteps in which we abstain, we can see that:
$\sum_{t\in[T]}\sum_{j\in[M]}\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket r_{t,b_{j}}-\sum_{t\in[T]}\mathbb{E}[r_{t,a_{t}}]=\sum_{t\in T^{\prime}}r_{t,b_{j}}-\sum_{t\in T^{\prime}}\mathbb{E}[r_{t,a_{t}}]\,,$
For any ball $j\in[M]$, we define $T_{j}=\{t\in[T^{\prime}]\,|\,\llbracket x_{t}\in\mathcal{B}_{j}\rrbracket\}$. Following the ideas of Seldin and Lugosi [2016], for any of the sets $T_{j}$ we can create a multi-armed bandit instance as the one described in the lower bound by Auer et al. [2002]. Note that in the lower bound construction, the abstention arm would be a forehand known suboptimal arm, which results in a lower bound of the order $c\sqrt{(K-1)T}$, for the constant $c=\frac{\sqrt{2}-1}{\sqrt{32\ln(4/3)}}>0$. Since the presented context $x_{t}$ is chosen adversarially at each time step, we can ensure that each basis element is activated for $|T^{\prime}|/M$ time steps, obtaining:
$\sum_{j\in[M]}\left(\sum_{s\in T^{\prime}_{j}}r_{s,b_{j}}-\sum_{s\in T^{\prime}_{j}}\mathbb{E}[r_{s,a_{s}}]\right)$ $\geq\sum_{j\in[M]}c\sqrt{(K-1)|T^{\prime}_{j}|}$
$=\sum_{j\in[M]}c\sqrt{(K-1)|T^{\prime}|/M}$
$=c\sqrt{M(K-1)|T^{\prime}|}$
As we can choose $|T^{\prime}|$ to be any fraction of T, we end up with the desired lower bound of the order $\Omega(\sqrt{MKT})$, which matches, up to logarithmic factors, the cumulative reward bound presented in Theorem 5.3. ∎
Appendix D Overlapping balls extension
@@
In this section, we present the theorem that allows us to present the results of overlapping balls as expressed in Section 5.2. Note that Theorem 5.3 is the special case of Theorem D.1 when the balls are disjoint and $u_{j}=1$ for all $j\in[M]$.
Theorem D.1.
@@
Let $M\in\mathbb{N}$ and $\{(B_{j},b_{j},u_{j})\,|\,j\in[M]\}$ be any sequence such that $B_{j}$ is a ball, $b_{j}\in[K]$ is an action, and $u_{j}\in[0,1]$ is such that for all $x\in\mathcal{X}$ we have:
$\sum_{j\in[M]}\llbracket x\in B_{j}\rrbracket u_{j}\leq 1\,.$
@@
${r}^{*}_{t}:=\sum_{j\in[M]}\llbracket x_{t}\in B_{j}\rrbracket u_{j}{r}_{t,b_{j}}\,,$
which represents the reward of the policy induced by $\{(B_{j},b_{j},u_{j})\,|\,j\in[M]\}$ on trial $t$. The regret of CBA, with the set of experts given in Section 5.2 and with correctly tuned parameters, is then bounded by:
@@
Direct from Theorem 3.1 using the experts (with efficient implementation) given in Section 5.2
Appendix E The details of the graph bases
This section expands the definition and explanations for the bases we used in the Experiment. Remember that we refer to any set of experts that correspond to set-action pairs of the form $(B,k)\in 2^{\mathcal{X}}\times[K]$ as a basis elements, and a set of basis elements as basis.
E.1 $p$-seminorm balls on graphs
As we see in Sec. 5.2, the CBA seems to work only for vector data. However, in the following sections, we explore how our CBA algorithm can be applied to graph data by creating a ball structure over the graph.
@@ E.1 $p$-seminorm balls on graphs
which is a well-defined metric for $p\in[1,\infty)$ if the graph is connected and may be defined for $p=\infty$ by taking the appropriate limits. When $p=2$ this is the square root of the effective resistance circuit between nodes $i$ and $j$ which comes from interpreting the graph as an electric circuit where the edges are unit resistors and the denominator of Equation (16) is the power required to maintain a unit voltage difference between $u$ and $v$ [Doyle and Snell, 1984]. More generally, $d_{p}(i,j)^{p}$ is known as $p$-(effective) resistance [Herbster and Lever, 2009, Alamgir and von Luxburg, 2011, Saito and Herbster, 2023]. When $p\in\{1,2,\infty\}$ there are natural interpretation of the $p$-resistance. In the case of $p=1$, we have that the effective is equal to one over the number of edge-disjoint paths between $i$ and $j$ which is equivalently one over the minimal cut that separates $i$ from $j$. When $p=2$ it is the effective resistance as discussed above. And finally when $p=\infty$ we have that $d_{\infty}$ is the geodesic distance (shortest path) between $i$ and $j$. Note that, interestingly, there are at most $2N$ distinct balls for $d_{1}$; as opposed to the general bound $O(N^{2})$ on the number of metric balls. This follows since $d_{1}$ is an ultrametric. A nice feature of metric balls is that they are ordinal, i.e., we can take an increasing function of the distance and the distinct are unchanged. The time complexity for each ball is as follows. For $d_{1}$ ball, we compute every pair of distance in $\mathcal{O}(N^{3})$ using the Gomory-Hu tree [Gomory and Hu, 1961]. For $d_{2}$ ball, it is actually enough to compute the pseudoinverse of graph Laplacian once, which costs $\mathcal{O}(N^{3})$ [Doyle and Snell, 1984]. For $d_{\infty}$ ball, we can compute every pair of distance in $\mathcal{O}(N^{3})$ by Floyd–Warshall algorithm [Floyd, 1962].
E.2 Community detection bases
@@
We can compute a basis for a given graph $G=(V,E)$ using community detection algorithms. Community detection is one of the most well-studied operations for graphs, where the goal is to find a partition $\{C_{1},\dots,C_{q}\}$ of $V$ (i.e., $\bigcup_{i=1}^{q}C_{i}=V$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$) so that each $C_{i}$ is densely connected internally but sparsely connected to the rest of the graph [Fortunato, 2010]. There are many community detection algorithms, all of which can be used here, but the most popular algorithm is the Louvain method [Blondel et al., 2008]. We briefly describe how this algorithm works. The algorithm starts with an initial partition $\{\{v\}\mid v\in V\}$ and aggregates the clusters iteratively: For each $v\in V$, compute the gain when moving $v$ from its current cluster to its neighbors’ clusters and indeed move it to a cluster with the maximum gain (if the gain is positive). Note that the gain is evaluated using modularity, i.e., the most popular quality function for community detection [Newman and Girvan, 2004]. The algorithm repeats this process until no movement is possible. Then the algorithm aggregates each cluster to a single super node (with appropriate addition of self-loops and change of edge weights) and repeats the above process on the coarse graph as long as the coarse graph is updated. Finally, the algorithm outputs the partition of $V$ in which each cluster corresponds to each super node in the latest coarse graph. Note that it is widely recognized that the Louvain method works in $\mathcal{O}(N\log N)$ in practice [Traag, 2015].
@@
To obtain a finer-grained basis, we apply the so-called greedy peeling algorithm for each $C_{i}$ in the output of the Louvain method. For $C_{i}\subseteq V$ and $v\in C_{i}$, we denote by $d_{C_{i}}(v)$ the degree of $v$ in the induced subgraph $G[C_{i}]$. For $G[C_{i}]$, the greedy peeling iteratively removes a node with the smallest degree in the currently remaining graph and obtains a sequence of node subsets from $C_{i}$ to a singleton. Specifically, it works as follows: Set $j\leftarrow|C_{i}|$ and $C_{i}^{(j)}\leftarrow C_{i}$. For each $j=|C_{i}|,\dots,2$, compute $v_{\text{min}}\in\operatorname*{arg\,min}\{d_{C_{i}^{(j)}}(v)\mid v\in C_{i}^{(j)}\}$ and $C_{i}^{(j-1)}\leftarrow C_{i}^{(j)}\setminus\{v_{\text{min}}\}$. Using a sophisticated data structure, this algorithm runs in linear time [Lanciano et al., 2024].
@@
In summary, our community detection basis is the collection of node subsets $\{C_{i}^{(j)}\mid i=1,\dots,q,\,j=1,\dots,|C_{i}|\}$ together with $\{\{v\}\mid v\in V\}$ for completeness.
E.3 Graph convexity bases